Check the continuity of the function


f: R^2to R at(0,0) , where f is defined by


f(x,y) = { 3x^2y/( x^2+y^2) , if (x,y)≠(0,0)

{ 3 , if (x,y)= (0,0)


1
Expert's answer
2021-06-21T06:09:05-0400

"If \space lim_{(x,y)\to(0,0)}f(x,y)=f(0,0)\\
\text{from any path then f is continuous.}\\
lim_{(x,y)\to(0,0)}f(x,y)=lim_{(x,y)\to(0,0)}\frac{3x^2y}{x^2+y^2}\space ( put \space y=mx)\\
=lim_{x\to0}\frac{3mx^3}{x^2(1+m^2)}\\
=lim_{x\to0}\frac{3mx}{1+m^2}\\
=0\neq f(0,0)=3\\
\text{Therefore, the given function is not continuous at (0,0).}"